20 mL of the standard hard water (containing 15g of CaCO3 per litre) required 10 mL of EDTA solution for endpoint. 100 mL of the water sample required 18 mL of EDTA solution. While the same volume of water after boiling consumed 12 mL of the EDTA solution. Calculate the carbonate and non-carbonate hardness present in water. Which buffer is used for the titration and at what pH?


1
Expert's answer
2021-03-16T07:52:54-0400

The standard water contains 15 g of CaCO3/l. 

1000 ml of standard hard water (SHW), therefore, contain 15,000 mg of CaCO3. 

1 ml of standard water = 15,000/1000 = 15 mg of CaCO3 equivalent. 

25 ml of EDTA solution = 20 ml of SHW contains of CaCO3 equivalent. 

1 ml of EDTA solution = 300/25 = 12 mg of CaCO3 equivalent. 

Calculation of hardness of water: 

100 ml of water 18 ml of EDTA solution 

 mg of CaCO3 equivalent. 

 1000 ml of water 2160 mg of CaCO3 equivalent. 

Hence, hardness of water = 2160 mg/l or ppm. 

Calculation of permanent hardness of water: 

100 ml of water (after boiling) 12 ml of EDTA solution 

 mg of CaCO3 equivalent. 

 1000 ml of water 1440 mg of CaCO3 equivalent. 

Hence, permanent hardness of water = 1440 ppm. 

Calculation of temporary hardness of water: 

Temporary hardness of water = (Total – Permanent) hardness 

Thus, 

Temporary hardness of water = 2160 – 1440 = 720 ppm. 

Answer: Temporary hardness of water = 720 ppm.